# DiaBlo: Diagonal Blocks Are Sufficient For Finetuning

### v1

@@ A.1 Commonsense Reasoning
The results in Table 1 on commonsense reasoning in FP16 follow the standard setting in [23, 33]. The results in Table 2 on commonsense reasoning in BF16 follow the setting in [61]. The hyper-parameteres for FP16 are provided in 7 and the hyper-parameteres for FP16 are provided in 8
@@
A.2 Arithmetic Reasoning
We follow the training setting of MiLoRA [49] for the arithmetic reasoning task. The hyper-parameter setting for fine-tuning Llama2-7B on MetaMathQA dataset is shown in Table 9.
@@ A.2 Arithmetic Reasoning
A.3 Code Generation and Safety Alignment
We follow the training setting of LoRI [61] for the code generation and safety alignment tasks. The hyper-parameter setting for code generation task is shown in Table 10 and the setting for the safety alignment is show in Table 11.
@@ A.3 Code Generation and Safety Alignment
LR scheduler linear constant linear
@@
A.4 Quantized Models
We finetune with quantized models on MATH-10K dataset. The fine-tuning hyper-parameters are show in Table 12.

### v2

@@ A.4.2 Commonsense Reasoning
The results in Table 1 on commonsense reasoning in FP16 follow the standard setting in (Hu et al., 2023; Liu et al., 2024). The results in Table 6 on commonsense reasoning in BF16 follow the setting in (Zhang et al., 2025a). The hyper-parameteres for FP16 are provided in Table 12 and the hyper-parameteres for BF16 are provided in Table 13.
@@
A.4.3 Arithmetic Reasoning
We follow the training setting of MiLoRA (Wang et al., 2024a) for the arithmetic reasoning task. The hyper-parameter setting for fine-tuning Llama2-7B on MetaMathQA dataset is shown in Table 14.
@@ A.4.3 Arithmetic Reasoning
A.4.4 Code Generation and Safety Alignment
We follow the training setting of LoRI (Zhang et al., 2025a) for the code generation and safety alignment tasks. The hyper-parameter setting for code generation task is shown in Table 15 and the setting for the safety alignment is show in Table 16.
@@ A.4.4 Code Generation and Safety Alignment
LR scheduler linear constant
@@
A.4.5 Quantized Models
We finetune with quantized models on MATH-10K dataset. The fine-tuning hyper-parameters are shown in Table 17.
@@ A.4.5 Quantized Models
A.4.6 MT-Bench
We finetune Llama2-7B on a 52k subset of WizardLM and evaluate on MT-Bench. The hyper-parameters are shown in Table 18.
Table 18: Hyper-parameters for the finetuning of Llama2-7B for MT-Bench evaluation.
Hyper-parameter Llama2-7B
number of blocks 320
Optimizer AdamW
Learning rate 1.5e-3
Weight decay 0
LR scheduler linear
Warmup steps 100
Epochs 1
Batch size 32
Max sequence length 1024
Modules Q K V O G U D
A.4.7 GLUE
We finetune the T5-base model on training datasets of GLUE tasks and evaluate on the development sets. The hyper-parameters are shown in Table 19.
Table 19: Hyper-parameters for the finetuning of T5-base on GLUE.
Hyper-parameter MRPC COLA SST-2 QNLI MNLI
number of blocks 64
Optimizer AdamW
Learning rate 1.4e-2 1e-2 5e-3 4e-3 2e-3
Weight decay 0
LR scheduler cosine
Warmup ratio 3%
Epochs 1
Batch size 32
Max sequence length 128
Modules All linear except embeddings and LM head
A.4.8 Retaining and fine-tuning sparse updates
For the results in Table 5, retaining sparse updates uses scales $s=0.75,0.75,0.7$ for sparsity $1/32,1/64,1/128$, respectively. The finetuning follows the setting in Table 20.
Table 20: Hyper-parameters for the finetuning of Llama3-8B on GSM8K.
Hyper-parameter Llama3-8B
Optimizer AdamW
Learning rate 4e-4
Weight decay 0
LR scheduler linear
Warmup steps 20
Epochs 2
Batch size 32
Max sequence length 512
Modules Q K V O G U D
A.5 The Use of Large Language Models
Parts of the manuscript were refined using large language model tools to improve clarity, grammar, and readability.